OpenStudy (anonymous):

What is the radius of the circular path? ANY HELP(: A proton moves at a speed of 8.4x10^4 m/s as it passes through a magnetic field of 12 mT.

OpenStudy (schrodingers_cat):

Well start off with the force on a moving charge through a magnetic field. F=q(vxB) (V is crossed With B) Using the definition of a cross product this reduces to F = QvBSin(theta) Assuming the magnetic filed is perpendicular with the velocity the sine of angle between them is 1. Equating this with the centripetal force for circular motion yields mv2/r=qvB Solving for r you get r=(mv)/(qb) From there just plug in the numbers. Hope this Helps :)

OpenStudy (schrodingers_cat):

There is a typo where i equated the forces it should be m(v^2)/r :)

OpenStudy (anonymous):

What should go where when plugging in? >.<

OpenStudy (anonymous):

Where; m= mass of the proton B= Magnetic field v= velocity of the proton